Find the relative extrema using both First and Second Partial Derivative test
fx=1+8x-3x2
1
Expert's answer
2020-12-29T15:22:05-0500

"f(x)=1+8x-3x^2"

Differentiating w.r.t "x" ,

"f'(x)=\frac{d}{dx}(f(x))=8-6x"

Differentiating w.r.t "x" ,

"f''(x)=\frac{d^2f(x)}{dx}=\frac{d}{dx}(f'(x))=-6"


First Derivative test :-

"f'(x)=0 \Rightarrow 8-6x=0 \Rightarrow x= \frac{4}{3}"

"\Rightarrow f(x)" has a critical point at "x=\frac{4}{3}"

If "x<\frac{4}{3}, f'(x)=8-6x>0"

"\Rightarrow" "f(x)" is increasing "\forall x<\frac{4}{3}"


If "x>\frac{4}{3}", "f'(x)=8-6x<0"

"\Rightarrow f(x)" is decreasing "\forall x>\frac{4}{3}"


Second Derivative test :-

"f''(\frac{4}{3})=-6<0" , as "f''(x)=-6 \forall x\in \R"


Now, "f(4/3)=1+\frac{32}{3}-\frac{16}{3}=\frac{19}{3}"


So, from First and Second derivative test, we can conclude that:-


"f(x)" has a relative maximum, "\frac{19}{3}" at "x=\frac{4}{3}" .

and, "f" doesn't have a relative minimum.


Graph :-